OpenStudy (anonymous):

solve for n: n-(-45)=90

OpenStudy (anonymous):

n = 45

jimthompson5910 (jim_thompson5910):

n-(-45)=90 n+45=90 n=90-45 n=45

OpenStudy (anonymous):

thank you!!!

jimthompson5910 (jim_thompson5910):

np, glad to be of help